# Characterization of a radiation detector based on opaque water-based liquid scintillator

### v1

@@ 1 Introduction
The two main categories of methods for obtaining event topological information in a traditional scintillation detector are imaging the interior of the detector, usually by using multiple photosensors spread out around the outside of the detector volume [Suzuki2019TheExperiment, Li2021KamLAND-ZenAnalysis, Wagner2018AMeasurements, Askins2020Theia:Detector, Harvey2021DevelopmentApplications, Manfredi2020TheCamera, Ranucci2001Borexino, Chen2008TheExperiment, DoubleChooz, An2016NeutrinoJUNO], and physically segmenting the detector into voxels [Li2019AArrays, Sutanto2021SANDD:Scintillator, Weinfurther2015LowCamera, Manfredi2020TheCamera, Ayres2007TheReport, Aliaga2014DesignDetector].
@@
Section 2 describes the experimental apparatus and materials, including the detector design, the data acquisition scheme, and the various scintillation liquids investigated. While several categories of scintillation liquids have been considered for opaque scintillation systems, including emulsion-based (opaque formulations of water-based liquid scintillators), wax-based (NoWaSH-20 [Buck2019]), and liquid scintillator cooled to liquid xenon temperatures [PandaX], in this study we focus on emulsion-based liquids. Section 3 contains the experimental results, including the measurement of light yields of each of the liquid scintillators and the determination of optical parameters for one of them. Then, response to gamma rays is shown for each scintillation liquid and the vertex reconstruction capability of one of them is demonstrated using pulses of photons from a fiber-coupled laser. Lastly, in Sec. 4, we give context for the achieved precision of vertex reconstruction and discuss planned future work.
@@ 2 Apparatus and material
The efficiency of light collection, transport, and conversion to signal for the experimental apparatus that shares most components with the one in this work was characterized in Ref. [Wilhelm2023EvaluationFibers]. The type of fiber (Kuraray Y-11(200)), 64-channel multianode photomultiplier tube (MAPMT) (Hamamatsu H12700A), and digitizer (CAEN V1730) were the same for this experiment as in that previous work. An opaque formulation of water-based liquid scintillator (discussed in Sec. 2.3) was chosen due to its promising optical characteristics and material compatibility, and is referred to as opaque Water-based Liquid Scintillator (oWbLS).
@@ 2.1 Detector design and construction
Figure 2 shows the experimental apparatus. The body of the detector was made from an additively manufactured material (VeroWhite), which is similar to acrylic. The detector volume was a 1 L cube, 10 cm on a side. There were 16 WLS fibers arranged in a grid in each vertical plane. The fibers in the $xz$ plane were offset 1 cm higher than the $yz$ plane to avoid collision and provide additional information for the $z$ direction. Figure 1 shows a sketch of the detector geometry. The fibers were glued in place with a fast-curing epoxy (5-minute Z-POXY), which was found to be compatible with WLS fibers and to cause less light loss than other glues tested [Bannister2023InvestigatingFibres].
Figure 1: A sketch of the detector geometry. The center panel shows a three-dimensional view of the detector body. The left panel shows the layout of the fibers in the $xz$ plane, and the right panel shows the layout of the fibers in the $yz$ plane.
@@ 2.2 WLS fibers and MAPMT
The WLS fibers were Kuraray Y-11(200) (non S-type) blue-to-green shifters, 1 mm in diameter [2022KurarayDatasheet]. The fibers were cut with a precision cleaver and then polished with fine-grit sandpaper. Both ends of each fiber terminated at a single pixel of the Hamamatsu H12700A MAPMT and were held in place by a printed interface and silicone optical grease. The pixel size of the MAPMT is 36 mm2, so the two fiber ends fit comfortably in the center of the pixel. The channels were chosen in a checkerboard pattern on the MAPMT to minimize crosstalk, as shown in Fig. 2(b).
@@ 2.3 Opaque water-based liquid scintillator
In traditional WbLS, water is used to reduce the light yield, whereas in oWbLS the water introduces opacity. The micelles constitute scattering centers, and the appearance of the liquid becomes like soapy water. The organic component of the WbLS used in this study is a mix of modified polyethylene glycol based surfactants in $>$90% of di-isopropylnaphthalene (DIN) liquid scintillator loaded with a fluor of 2,5-diphenyloxazole (PPO) at 3g/L. In addition, the wavelength shifter 1,4-Bis(2-methylstyryl)benzene (bis-MSB) at 15mg/L was included to match the emission spectrum with the absorption spectrum of the WLS fibers.
Images of three formulations of oWbLS (oWbLS 1–3) are shown in Fig. 3.
@@ 3 Results
The light yield of oWbLS was measured using the method described in Ref. [Cumming2019ImprovingScintillators], in which a large photomultiplier tube views a cuvette of liquid scintillator exposed to gamma rays from 137Cs. The light yield of an unknown scintillator is determined by comparison to a characterized reference. The reference scintillator was linear alkylbenzene (LAB) with 3 g/L of PPO, which has a light yield between 8700 [Buck2019ProductionExperiment] and 11100 [Caravaca2020CharacterizationSeparation] photons/MeV. The average of these two values was taken as the reference value for the light yield of LAB+PPO. The light yield of oWbLS was determined to be 12000$\pm$2000 photons/MeV. This value is consistent with other scintillators based on DIN, such as EJ-309, which has a light yield of 12300 photons/MeV [EJ-309Technology]. The uncertainty is dominated by the range of light yields reported in the literature for the reference scintillator. These results are summarized in Table 1.
@@
First, not all photons generated in this prototype will reach a WLS fiber. This could be improved in future designs by increasing the absorption length of the scintillator and by optimizing the fiber pitch for a given set of scintillator optical properties. Second, only some of the photons that reach a fiber are absorbed and re-emitted. This could be optimized through a better matching of the WLS absorption spectrum and the scintillator emission spectrum. Third, most wavelength-shifted photons are re-emitted in a direction that does not result in total internal reflection and thus are not transported to the photosensor. Instead, they immediately escape the fiber. This causes the greatest loss of any of the steps described here. Kuraray estimates that their multi-clad fibers have a trapping efficiency of at least 10.8% [2022KurarayDatasheet], meaning almost 90% of photons that are emitted in the core of a WLS fiber are lost. These photons have already been wavelength-shifted, and therefore cannot be reabsorbed by another WLS fiber. Investigation into the improvement of the effective fiber trapping efficiency is ongoing. Fourth, some photons that are transported in the fiber core are absorbed before reaching the photosensor. Because all the fibers in this design terminate at a single MAPMT, the fibers are much longer than they would need to be if the photosensors were located immediately outside the active volume. Last, less than half of the photons that reach the photosensor will generate a measurable electric signal, a probability described as the device quantum efficiency. This loss is not unique to LiqudO detectors, but is unavoidable in any detector that uses light to detect radiation. However, the quantum efficiency of the MAPMT used in this prototype is less than a typical SiPM, so future designs have the opportunity for improvement in this step as well.
@@
The degree of light confinement in an opaque scintillator depends on its scattering and absorption lengths, as well as the anisotropy of the scattering. In this analysis, we account for the effect of scattering anisotropy by using the reduced scattering coefficient
@@
Figure 5: Relationship between the number of triggered channels and total event signal for the three formulations of oWbLS. The data shown was generated by exposing the detector to gamma rays from 60Co. A liquid scintillator with an extremely short scattering length (e.g. 1% of the fiber pitch) would produce events in which only one or two channels were triggered, regardless of the total intensity of the event. Thus, the more channels triggered per event for a given summed pulse integral value, the less confining the liquid was to the scintillation photons.
@@
The measured distribution of signal over WLS fiber channels was compared to a Geant4 simulation in which six parameters were allowed to vary: the $x$, $y$, and $z$ position of the fiber tip, the absorption and reduced scattering length of the oWbLS, and the reflectivity of the detector walls. By optimizing these parameters to minimize the residual between simulated and experimental data, estimates of the parameters can be made. The estimated position of the fiber tip before fitting the data was $(0\pm 5,0\pm 5,0\pm 5)$ mm, with the center of the detector volume taken as the origin.
The reflectivity of the walls was estimated from Ref. [Golhin2023AppearanceJetting], which quantified the optical parameters of several colors of printed plastics. The wavelength of the laser ($407$ nm) lies in the transition region for the spectral reflectance of VeroWhite plastics. The indices of refraction of liquid scintillation cocktails based on DIN were found to be between 1.55 and 1.6 at 404.7 nm, with most around 1.56 [Kossert2013MeasurementCocktails]. The index of refraction of oWbLS is likely to be slightly smaller as a result of the addition of water. The index of refraction of VeroClear, a transparent but similar material to VeroWhite, was reported to be about 1.48 at 400 nm [Song2015Matching-index-of-refractionVisualization], so the specular reflection at the oWbLS-VeroWhite boundary is expected to be less than in the measurement of Ref. [Golhin2023AppearanceJetting], which was done in air.
The optimal parameter combination was found with a genetic algorithm. In a genetic algorithm, different combinations of parameters represent individuals in a population. Each individual can be assessed for fitness, and the average fitness of a population can be made to increase over many generations. For each individual in a generation, a Geant4 simulation was run with the corresponding parameter values. The output of the simulation was the number of photoelectrons per channel per pulse, just as in the experiment. The fitness of an individual was calculated as the value of $\chi^{2}$ between the simulated and experimental data. A comparison of the experimental and simulated signal distributions is shown in Fig. 7, and the set of optimized parameters is shown in Tab. 2. The number of photoelectrons detected in each channel per pulse depended on the proximity of the WLS fiber to the laser-coupled fiber.
@@
The absorption length of DIN is known to be shorter than LAB, but DIN could nevertheless be desirable for its higher light yield. The measured absorption lengths of pure DIN at 430 nm range from 1 m [Bonhomme2022SafeDetectors] to 4.2 m [Song2013FeasibilityScintillator] and are expected to be lower at shorter wavelengths. The value of 0.169 m found here is therefore lower than expected. There are several possible explanations for this discrepancy. First, the light attenuation during transport in WLS fibers may be greater than modeled. While the model of light transport in the fibers used here was previously validated [Wilhelm2023EvaluationFibers], there are several differences in this experimental campaign from prior experiments. One is that the fibers were occasionally exposed to ambient fluorescent room lighting for a period of about a year, and may have degraded. Secondly, while the fiber bending radius is within the manufacturer’s recommendation, the bending necessary to route the fibers to the MAPMT may nevertheless introduce unaccounted losses. The oWbLS is not pure DIN and may have a shorter absorption length due to the addition of the surfactant. Furthermore, the optical properties of the oWbLS may have degraded over time due to exposure to ambient oxygen and light. While the absorption length calculated here primarily served to enable simulations for vertex reconstruction in Sec. 3.5, it may not accurately reflect the true optical properties of the material. Additional experiments are planned to directly measure the optical properties of oWbLS and address these discrepancies.

### v2

@@ 1 Introduction
The two main categories of methods for obtaining event topological information in a traditional scintillation detector are imaging the interior of the detector, usually by using multiple photosensors spread out around the outside of the detector volume [Suzuki2019TheExperiment, Li2021KamLAND-ZenAnalysis, Wagner2018AMeasurements, Askins2020Theia:Detector, Harvey2021DevelopmentApplications, Manfredi2020TheCamera, Ranucci2001Borexino, Chen2008TheExperiment, DoubleChooz, An2016NeutrinoJUNO], and physically segmenting the detector into voxels [Li2019AArrays, Sutanto2021SANDD:Scintillator, Weinfurther2015LowCamera, Manfredi2020TheCamera, Ayres2007TheReport, Aliaga2014DesignDetector, Michael2008TheExperiment, Abreu2021SoLid:Experiment, Ashenfelter2019TheExperiment, Abe2011TheExperiment, Blondel2018AViews].
@@
Section 2 describes the experimental apparatus and materials, including the detector design, the data acquisition scheme, and the various scintillation liquids investigated. While other categories of scintillators have been considered for opaque scintillation detectors, including wax-based (NoWaSH-20 [Buck2019]) and liquid scintillator cooled to liquid xenon temperatures [PandaX], in this study we introduce opaque emulsion-based liquid scintillators based on the principle of water-based liquid scintillator. Section 3 contains the experimental results, including the measurement of light yields of each of the liquid scintillators and the determination of optical parameters for one of them. Then, response to gamma rays is shown for each scintillation liquid and the vertex reconstruction capability of one of them is demonstrated using pulses of photons from a fiber-coupled laser. Lastly, in Sec. 4, we give context for the achieved precision of vertex reconstruction and discuss planned future work.
@@ 2 Apparatus and material
The efficiency of light collection, transport, and conversion to signal for an experimental apparatus that shares most components with the one in this work was characterized in Ref. [Wilhelm2023EvaluationFibers]. The type of fiber (Kuraray Y-11(200)), 64-channel multianode photomultiplier tube (MAPMT) (Hamamatsu H12700A), and digitizer (CAEN V1730) were the same for this experiment as in that previous work. An opaque formulation of water-based liquid scintillator (discussed in Sec. 2.3) was chosen due to its promising optical characteristics and material compatibility, and is referred to as opaque Water-based Liquid Scintillator (oWbLS). [^1] Water-based liquid scintillator (WbLS) is often, but not always, majority water by volume. This term clarifies that water is added to a traditional scintillator (oil-based solvent) via stable emulsion to create a different type of scintillator. The formulation described in this experiment includes a small fraction of water to promote scattering, and the majority of its volume is an oil-based solvent.
@@ 2.1 Detector design and construction
Figure 1 shows the experimental apparatus. The body of the detector was made from an additively manufactured material (VeroWhite). The wall reflectivity is discussed in Sec. 3.2. Due to the inherent light confinement of the design and the establishment of a fiducial region away from the walls, reflected light plays an insignificant role in the reconstructions of point-like events (further discussed in Sec. 3.5). The detector volume was a 1 L cube, 10 cm on a side. There were 16 WLS fibers arranged in a grid in each vertical plane. The fibers in the $xz$ plane were offset 1 cm higher than the $yz$ plane to avoid collision and provide additional information for the $z$ direction. Figure 2 shows a sketch of the detector geometry. The fibers were glued in place with a fast-curing epoxy (5-minute Z-POXY), which was found to be compatible with WLS fibers and to cause less light loss than other glues tested [Bannister2023InvestigatingFibres].
@@
Figure 2: A sketch of the detector geometry. (a) The layout of the fibers in the $xz$ plane, (b) a three-dimensional view of the detector body, and (c) the layout of the fibers in the $yz$ plane. The origin of the coordinate system is the center of the detector volume.
@@ 2.2 WLS fibers and MAPMT
The WLS fibers were Kuraray Y-11(200) (non S-type) blue-to-green shifters, 1 mm in diameter [2022KurarayDatasheet]. The fibers were cut with a precision cleaver and then polished with fine-grit sandpaper. Both ends of each fiber terminated at a single pixel of the Hamamatsu H12700A MAPMT and were held in place by a printed interface and silicone optical grease. The pixel size of the MAPMT is 36 mm2, so the two fiber ends fit comfortably in the center of the pixel. The channels were chosen in a checkerboard pattern on the MAPMT to minimize crosstalk, as shown in Fig. 1(b). The quantum efficiency of the MAPMT peaks at $\sim$32at 330 nm, but is $\sim$16at 500 nm, which is the approximate emission wavelength of Y-11 fibers [2022KurarayDatasheet, 2019HamamatsuDatasheet].
@@ 2.3 Opaque water-based liquid scintillator
In traditional WbLS, water is used to reduce the light yield, whereas in oWbLS the water introduces opacity. The micelles constitute scattering centers, and the appearance of the liquid becomes like soapy water. The organic component of the WbLS used in this study is a mix of modified polyethylene glycol-based surfactants in $>$90% of di-isopropylnaphthalene (DIN) liquid scintillator loaded with a fluor of 2,5-diphenyloxazole (PPO) at 3 g/L. We made three oWbLS formulations, the difference between which is the water content in the organic mixture. Different water content induced different opacities (scattering lengths) in the oWbLS. In this study, 1.5% of water was introduced in oWbLS1 while 2% and 2.5% water were added to the oWbLS2 and oWbLS3, respectively, to study detector response with different liquid scattering lengths. In addition, the wavelength shifter 1,4-Bis(2-methylstyryl)benzene (bis-MSB) was included at 15mg/L to match the emission spectrum with the absorption spectrum of the WLS fibers.
Images of the three formulations of oWbLS (oWbLS 1–3) are shown in Fig. 3.
@@ 3 Results
The light yield of oWbLS was measured using the method described in Ref. [Cumming2019ImprovingScintillators], in which two large photomultiplier tubes view a cuvette of liquid scintillator exposed to gamma rays from 137Cs. The apparatus was a Beckman LS-6500 liquid scintillation counter. The gamma rays from 137Cs interact with the scintillator samples primarily through Compton scattering, which results in a distribution of pulse heights that includes a discernible Compton edge. The technique described in Ref. [Cumming2019ImprovingScintillators] is a way to extract the location of the Compton edge. It involves first converting the light output spectrum to a power spectrum, fitting the power spectrum with a function, and then taking the derivative of that function. The location of the Compton edge is taken to be the maximum of the derivative of the fitted function. The light yield of an unknown scintillator is determined by comparing the location of the Compton edge of the unknown sample to a characterized reference. The reference scintillator was linear alkylbenzene (LAB) with 3 g/L of PPO, which has a light yield between 8700 [Buck2019ProductionExperiment] and 11100 [Caravaca2020CharacterizationSeparation] photons/MeV. The average of these two values was taken as the reference value for the light yield of LAB+PPO. The light yield of oWbLS was determined to be 12000$\pm$2000 photons/MeV. This value is consistent with other scintillators based on DIN, such as EJ-309, which has a light yield of 12300 photons/MeV [EJ-309Technology]. The uncertainty is dominated by the range of light yields reported in the literature for the reference scintillator. These results are summarized in Table 1.
@@
First, not all photons generated in this prototype will reach a WLS fiber. This could be improved in future designs by increasing the absorption length of the scintillator and by optimizing the fiber pitch for a given set of scintillator optical properties. Second, only some of the photons that reach a fiber are absorbed and re-emitted. This could be optimized through better matching the WLS absorption spectrum to the scintillator emission spectrum. Third, most wavelength-shifted photons are re-emitted in a direction that does not result in total internal reflection and thus are not transported to the photosensor. Instead, they immediately escape the fiber. This causes the greatest loss of any of the steps described here. Kuraray estimates that their multi-clad fibers have a trapping efficiency of at least 10.8% for double-end fiber readout [2022KurarayDatasheet], meaning that almost 90% of photons emitted in the WLS fiber core are lost. These photons have already been wavelength-shifted, and therefore cannot be reabsorbed by another WLS fiber. Investigation into the improvement of the effective fiber trapping efficiency is ongoing. Fourth, some photons that are transported in the fiber core are absorbed before reaching the photosensor. Because all the fibers in this design terminate at a single MAPMT, the fibers are much longer than they would need to be if the photosensors were located immediately outside the active volume. Last, less than half of the photons that reach the photosensor will generate a measurable electric signal, a probability described as the device quantum efficiency. This loss is not unique to LiqudO detectors but is unavoidable in any detector that uses light to detect radiation. However, the quantum efficiency of the MAPMT used in this prototype is less than a typical SiPM, so future designs have the opportunity for improvement in this step as well.
@@
The degree of light confinement in an opaque scintillator depends on its scattering and absorption lengths, as well as the scattering anisotropy. In this analysis, we account for the effect of scattering anisotropy by using the reduced scattering coefficient
@@
The Geant4 model uses Rayleigh scattering, which is isotropic [Saidi1995MieSkin]. In reality, the scattering may be in the Mie regime, meaning that the scattering center size is comparable to light wavelength, and is thus anisotropic [Bunge2006RayleighFibers]. The use of reduced scattering length in the model accounts for the potential anisotropy of Mie scattering in the experiment.
@@
Figure 5: Relationship between the number of triggered channels and total event signal for the three formulations of oWbLS. The data shown was generated by exposing the detector to gamma rays from a 60Co source placed on the center of the detector lid. 60Co emits two gamma rays, at 1173.2 keV and 1332.5 keV. The maximum energy deposition from a single scatter of the 1332.5 keV gamma ray is 1118 keV. A liquid scintillator with an extremely short scattering length (e.g. 1% of the fiber pitch) would produce events in which only one or two channels were triggered, regardless of the total intensity of the event. The more channels triggered per event for a given summed pulse integral, the less confining the liquid is to the scintillation photons.
@@
The measured distribution of signal over WLS fiber channels was compared to a Geant4 simulation. The model includes tracking each photon until it is ultimately absorbed, escapes the detector volume, or reaches the photocathode of the MAPMT. When a photon reaches the photocathode of the MAPMT, an application of the wavelength-dependent quantum efficiency of the MAPMT dictates whether it is counted as a hit. The model of WLS fiber includes both layers of the cladding and the wavelength-shifting process that takes place in the core. The parameters used to model fibers, including their absorption and emission spectra, are described in detail in Ref. [Wilhelm2023EvaluationFibers]. Six parameters were allowed to vary: $x$, $y$, and $z$ position of the fiber tip, absorption and reduced scattering length of the oWbLS, and the reflectivity of the detector walls. The photons in the simulation were given an initial momentum in the $-z$ direction to model the photons emitted from the laser fiber. By optimizing these parameters to minimize the residual between simulated and experimental data, estimates of the parameters can be made. The estimated position of the fiber tip before fitting the data was ($0\pm 5$(x),$0\pm 5$(y),$0\pm 5$(z)) mm, with the center of the detector volume taken as the origin.
The wall reflectivity was estimated from Ref. [Golhin2023AppearanceJetting], which quantified the optical parameters of several colors of printed plastics, to be approximately 0.3. The wavelength of the laser ($407$ nm) lies in the transition region for the spectral reflectance of VeroWhite plastics. The indices of refraction of liquid scintillation cocktails based on DIN were found to be between 1.55 and 1.6 at 404.7 nm, with most around 1.56 [Kossert2013MeasurementCocktails]. The index of refraction of oWbLS is likely to be slightly smaller as a result of the addition of water. The index of refraction of VeroClear, a transparent but similar material to VeroWhite, was reported to be about 1.48 at 400 nm [Song2015Matching-index-of-refractionVisualization], so the specular reflection at the oWbLS-VeroWhite boundary is expected to be less than in the measurement of Ref. [Golhin2023AppearanceJetting], which was done in air.
The optimal parameter combination was found with a genetic algorithm, using the Multi-variate Analysis (TMVA) package in ROOT. In a genetic algorithm, different combinations of parameters represent individuals in a population. Each individual can be assessed for fitness, and the average fitness of a population can be made to increase over many generations. For each individual in a generation, a Geant4 simulation was run with the corresponding parameter values. The output of the simulation was the number of photoelectrons per channel per pulse, just as in the experiment. The fitness of an individual was calculated as the value of $\chi^{2}$ between the simulated and experimental data. A comparison of the experimental and simulated signal distributions is shown in Fig. 7, and the set of optimized parameters is shown in Table 2. The experimental data overlaid onto the two detector planes are shown in Fig. 8. The number of photoelectrons detected in each channel per pulse depends on the proximity of the WLS fiber to the laser-coupled fiber. Since the light confinement is stochastic, it is possible for photons to propagate to more distant fibers. This is apparent by the flat yield of $<1$ photons/pulse for fibers in the corners of the detector shown in Fig. 8.
@@
The absorption length of DIN is known to be shorter than LAB, but DIN could nevertheless be desirable for its higher light yield. The measured absorption lengths of pure DIN at 430 nm range from 1 m [Bonhomme2022SafeDetectors] to 4.2 m [Song2013FeasibilityScintillator] and are expected to be lower at shorter wavelengths. The value of 0.169 m found here is therefore lower than expected. There are several possible explanations for this discrepancy. First, the light attenuation during transport in WLS fibers may be greater than modeled. While the model of light transport in the fibers used here was previously validated [Wilhelm2023EvaluationFibers], there are several differences in this experimental campaign from prior experiments. One is that the fibers were occasionally exposed to ambient fluorescent room lighting for about a year, and may have degraded. A dedicated study on the effect of ambient light on WLS fibers has not been carried out, but Ref. [Chung1993EffectsFibers] found that a permanent 38% reduction of the attenuation length in a single-clad Kuraray scintillating fiber could be induced by a 137-hour exposure to ambient light. Kuraray estimates the attenuation length of Y-11 fiber to be 3.5 m [2022KurarayDatasheet], which becomes 2.2 m after a 38% degradation. Assuming a 0.5 m path for a photon to reach the MAPMT, this would result in an 8% signal loss in our detector. Secondly, while the fiber bending radius is within the manufacturer’s recommendation, the bending necessary to route the fibers to the MAPMT may nevertheless introduce unaccounted losses. The oWbLS is not pure DIN and may have a shorter absorption length due to the addition of the surfactant. Furthermore, the optical properties of the oWbLS may have degraded over time due to exposure to ambient oxygen and light. While the absorption length calculated here primarily served to enable simulations for vertex reconstruction in Sec. 3.5, it may not accurately reflect the true optical properties of the material. Additional experiments, described in Sec. 4.2, are planned to directly measure the optical properties of oWbLS and address these discrepancies.
@@
Figure 8: The experimental data from Fig. 7 shown overlaid on a detector sketch. The tip of the arrow shows the terminus position of the laser fiber. The light ball is located approximately one mean free path directly below the laser fiber tip.